# Understanding and Improving Shampoo and SOAP via Kullback–Leibler Minimization

### v2

@@ Abstract
Shampoo and its efficient variant, SOAP, use structured second-moment estimation and have attracted growing interest for their effectiveness in training neural networks (NNs). In practice, Shampoo requires step-size grafting with Adam to achieve competitive performance. SOAP mitigates this by applying Adam in Shampoo’s eigenbasis and further reducing per-iteration runtime. However, reliance on Adam introduces additional memory overhead in both methods. Prior theoretical interpretations have primarily examined their estimation schemes using the Frobenius norm. Motivated by the natural correspondence between the second moment and a covariance matrix, we reinterpret the estimation procedures in Shampoo and SOAP as instances of covariance estimation through the lens of Kullback-Leibler (KL) divergence minimization. This perspective reveals a previously overlooked theoretical limitation and motivates principled improvements to their design. Building on the KL perspective, we propose practical estimation schemes—KL-Shampoo and KL-SOAP—that match or exceed the performance of Shampoo and SOAP for pre-training various NNs while maintaining SOAP-level per-iteration runtime. Notably, KL-Shampoo does not rely on Adam to achieve superior performance, thereby avoiding the associated memory overhead. Surprisingly, KL-Shampoo consistently outperforms the other methods in our experiments.
@@ 1 Introduction
Optimizer Shampoo has recently dethroned Adam (Kingma & Ba, 2015) as the winner of several competitions in training a wide range of neural network (NN) models (Dahl et al., 2023; Kasimbeg et al., 2025). Consequently, Shampoo (Gupta et al., 2018) and its variant, SOAP (Vyas et al., 2025a), have drawn increasing attention. In practice, Shampoo requires step-size grafting with Adam to achieve competitive performance (Agarwal et al., 2020; Shi et al., 2023). SOAP addresses this by applying Adam in Shampoo’s eigenbasis and further reducing per-iteration runtime. However, reliance on Adam introduces additional memory overhead in both methods. Prior work (Morwani et al., 2025; Eschenhagen et al., 2025; An et al., 2025; Xie et al., 2025) has investigated their structural preconditioner schemes—which approximate the flattened gradient 2nd moment (Duchi et al., 2011)—through the Frobenius norm. However, few studies have examined these schemes from the perspective of Kullback–Leibler (KL) divergence. Compared to the Frobenius norm, the KL divergence between zero-mean Gaussian covariance matrices is more appropriate for interpreting Shampoo’s and SOAP’s preconditioners as Gaussian covariance estimators (Amari, 2016; Minh & Murino, 2017), since the second moment they approximate can be viewed as the covariance matrix of a zero-mean Gaussian. A similar KL perspective has provided a unified framework to interpret (Fletcher, 1991; Waldrip & Niven, 2016) and extend (Kanamori & Ohara, 2013a, b) structural preconditioner estimation in quasi-Newton methods such as BFGS and DFP—something the Frobenius norm does not. Moreover, the KL divergence intrinsically respects the symmetric positive-definite constraint (Amari, 2016; Minh & Murino, 2017) that preconditioners in Shampoo and SOAP must satisfy as adaptive (preconditioned) methods (Nesterov et al., 2018)—a property the Frobenius norm lacks. This constraint implies that the entries of the preconditioning matrix do not play equivalent roles and therefore should not be treated equally (Pennec et al., 2006; Bhatia, 2007)—a point the Frobenius norm ignores.
@@
In this work, we introduce a KL perspective that interprets the estimation schemes of Shampoo and SOAP as solutions to KL-minimization problems for covariance estimation. Our approach naturally extends to tensor-valued settings, where some existing theoretical interpretations may not apply. This perspective reveals a key limitation (illustrated in Fig.˜5): the Kronecker-structured estimators used by Shampoo and SOAP do not adequately solve the corresponding KL-minimization problem. This limitation, in turn, opens new opportunities for improvement. Leveraging this insight, we refine the estimation rules of Shampoo and SOAP and develop practical KL-based schemes—KL-Shampoo and KL-SOAP—that meet or exceed the performance of Shampoo and SOAP for NN (pre-)training while maintaining SOAP-level per-iteration runtime. Notably, KL-Shampoo does not rely on Adam to achieve superior performance, thereby avoiding Adam’s additional memory overhead (Table˜1). Empirically (see Fig.˜1), we show that KL-based methods are competitive for training a range of NNs and remain as flexible as Shampoo and SOAP for tensor-valued weights. Surprisingly, KL-Shampoo consistently outperforms the other methods in our experiments.
@@ 2 Background
For presentation simplicity, we focus on matrix-valued weights and the optimization update for a single parameter matrix ${\bm{\Theta}}\in\mathcal{R}^{d_{a}\times d_{b}}$, rather than a set of weight matrices for NN training. We use $\operatorname{Mat}(\cdot)$ to unflatten its input vector into a matrix and $\operatorname{vec}(\cdot)$ to flatten its input matrix into a vector. For example, ${\bm{\theta}}:=\operatorname{vec}({\bm{\Theta}})$ is the flattened weight vector and ${\bm{\Theta}}\equiv\mathrm{Mat}({\bm{\theta}})$ is the original (unflattened) weight matrix. Vector $\bm{g}$} is a (flattened) gradient vector for the weight matrix. We denote $\gamma$, $\beta_{2}$ and $\bm{S}$} to be a step size, a weight for moving average, and a preconditioning matrix for an adaptive method, respectively. $\mathrm{Diag}(\cdot)$ returns a diagonal matrix whose diagonal entries are given by its input vector, whilst $\mathrm{diag}(\cdot)$ extracts the diagonal entries of its input matrix as a vector.
@@
Using Adam for Shampoo’s stabilization increases memory usage. If the eigendecomposition is applied infrequently to reduce iteration cost, Shampoo has to apply step-size grafting with Adam to maintain performance (Agarwal et al., 2020; Shi et al., 2023) as empirically shown in Fig.˜2. Unfortunately, this increases its memory usage introduced by Adam (see Table˜1).
@@
Runing Adam in the eigenbasis increases memory usage. Introducing Adam’s (RMSProp’s) 2nd moment estimation increases SOAP’s memory consumption (see Table˜1). This is because this estimation, ${\bm{d}}\in\mathcal{R}^{d_{a}d_{b}\times 1}$, uses extra memory and cannot be expressed as a Kronecker product of any two vectors, such as ${\bm{d}}\neq{\bm{r}}_{a}\otimes{\bm{r}}_{b}$, where ${\bm{r}}_{a}\in\mathcal{R}^{d_{a}\times 1}$ and ${\bm{r}}_{b}\in\mathcal{R}^{d_{b}\times 1}$.
@@
Table 1: Memory usage of each method considered in this work. We store and update all of them in half precision (bfloat16). The memory overhead introduced by Adam is highlighted in red. Note that SOAP’s and KL-SOAP’s preconditioners, ${\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$, can not be expressed as a Kronecker product due to the augmented eigenvalues $\bm{d}$}, while Shampoo’s and KL-Shampoo’s preconditioners, ${\bm{Q}}\mathrm{Diag}({\bm{\lambda}}_{a}\otimes{\bm{\lambda}}_{b}){\bm{Q}}^{\top}$, can, where ${\bm{Q}}:={\bm{Q}}_{a}\otimes{\bm{Q}}_{b}$.
@@ 3 Second Moment Estimation via Kullback–Leibler Minimization
Many works (Morwani et al., 2025; Eschenhagen et al., 2025; An et al., 2025; Xie et al., 2025) primarily focus on matrix-valued weights and interpret Shampoo’s and SOAP’s estimation rules from the Frobenius-norm perspective. However, this norm does not account for the SPD constraint implicitly imposed on Shampoo’s and SOAP’s preconditioners, which ensures that the preconditioned gradient direction is a descent direction (Nesterov et al., 2018). As emphasized in the literature (Pennec et al., 2006; Bhatia, 2007), it is more appropriate to consider a “distance” that respects this constraint. We adopt the KL divergence because it naturally incorporates the SPD constraint, is widely used for covariance estimation (Amari, 2016; Minh & Murino, 2017), and provides a unified framework for reinterpreting and improving Shampoo’s estimation—even in tensor-valued settings where existing interpretations based on singular value decomposition (Van Loan & Pitsianis, 1993) may not apply. In numerical optimization, the KL divergence, known as a merit function (Byrd & Nocedal, 1989), offers a unifying interpretation (Fletcher, 1991; Waldrip & Niven, 2016) and extension (Kanamori & Ohara, 2013a, b) of the low-rank estimation schemes of BFGS and DFP. In contrast, the standard Frobenius norm cannot recover these updates without additional weighting (see Sec. 6.1 of Nocedal & Wright (2006)). In statistical estimation and inference, this KL divergence is also preferred over the Frobenius norm (James et al., 1961; Kivinen & Warmuth, 1999; Davis & Dhillon, 2006; Khan & Lin, 2017; Lin et al., 2019; Kunstner et al., 2021).
@@ 3.2 Improving Shampoo’s estimation: Idealized KL-Shampoo
Our KL perspective reveals a key limitation—empirically demonstrated in Fig.˜5—of Shampoo’s Kronecker estimation with $p=\nicefrac{{1}}{{2}}$ as a one-sided approach: it does not adequately solve the KL-minimization problem when both factors are learned jointly. Motivated by this observation, we design an improved estimation rule that updates the two factors simultaneously. We refer to this scheme as idealized KL-Shampoo, which is a two-sided approach.
@@
˜2 (proof in Appx.˜B) establishes a closed-form condition (see Eq.˜4) when simultaneously learning both Kronecker factors to minimize the KL problem. Statistically, this condition corresponds to the maximum-likelihood covariance estimator of a zero-mean matrix Gaussian (Dutilleul, 1999). In machine learning, Lin et al. (2019, 2024) treated this condition as a theoretical example of a multilinear exponential-family (see Sec 5 of Lin et al. (2019)) for Kronecker-based optimization via natural gradient descent on an exponential-family manifold, while more recently, Vyas et al. (2025b) considered a similar condition motivated heuristically by gradient whitening. However, we cannot directly use this condition due to the correlated update between ${\bm{S}}_{a}$ and ${\bm{S}}_{b}$. For example, solving ${\bm{S}}_{a}^{*}$ requires knowing ${\bm{S}}_{b}^{*}$ in Eq.˜4 or vice versa. In practice, this optimal condition is often unachievable because the expectations in Eq.˜4 have to be approximated. Thus, we consider an estimated ${\bm{S}}_{k}$ to approximate ${\bm{S}}_{k}^{*}$ for $k\in\{a,b\}$ and propose an exponential moving average (EMA) scheme in practice:
@@
3.3 Comparison with Kronecker Schemes Using Alternative “Distances”
@@
Morwani et al. (2025) also consider a two-sided Shampoo variant based on the Frobenius norm and derive the optimal solution via rank-1 singular value decomposition (SVD) of the outer product $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ (Van Loan & Pitsianis, 1993). However, this optimal solution is often unattainable in practice and is computationally expensive for two reasons: (1) the expectation $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ must be approximated; and (2) performing the SVD itself is costly—yielding complexity ($O(d_{a}^{2}d_{b}^{2})$) even for rank-1 SVD—which is higher than the eigendecompositions with complexity ($O(d_{k}^{3})$) for $k\in\{a,b\}$ that we explicitly aim to avoid. Instead, we analyze the stationarity conditions (˜6, Appx.˜F) and derive a new variant, idealized F-Shampoo (Fig.˜6, Appx.˜F), that is structurally similar to KL-Shampoo. While a straightforward implementation of F-Shampoo performs poorly in practice, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted to improve its performance, as shown in Fig.˜7 (LABEL:{app:extra_exp}).
@@
Another variant, often discussed in the literature (Morwani et al., 2025; Vyas et al., 2025a; Eschenhagen et al., 2025), is Shampoo with trace scaling. Vyas et al. (2025a) established that Shampoo with trace scaling is equivalent to running Adafactor (Shazeer & Stern, 2018) in Shampoo’s eigenbasis. In contrast, KL-Shampoo is not equivalent to running Adafactor in its eigenbasis. To clarify this distinction, we make the theoretical connection between Shampoo and Adafactor more explicit: Shampoo with trace scaling is exactly a matrix generalization of Adafactor obtained by minimizing the von Neumann (VN) divergence (Tsuda et al., 2005; Dhillon & Tropp, 2008; Nock et al., 2012) and recovers Adafactor when its Kronecker factors are restricted to be diagonal, as we establish in ˜7 (Appx.˜G). By contrast, KL-Shampoo minimizes the KL divergence instead of the VN divergence. A straightforward implementation of Shampoo with trace scaling—referred to as idealized VN-Shampoo—performs poorly in practice. Howeve, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted (Fig.˜8, Appx.˜G) to substantially improve its performance, as shown in Fig.˜10 (LABEL:{app:extra_exp}).
@@
A natural question then arises: which “distance” is more suitable? Theoretically, the KL divergence is broadly applicable to arbitrary SPD matrices (Bhatia, 2007; Boumal et al., 2014) and is widely used for covariance matrices (Minh & Murino, 2017). In contrast, the Frobenius norm does not respect the SPD constraint, and the VN divergence is primarily motivated, studied, and applied for unit-trace SPD matrices (Tsuda et al., 2005; Nielsen & Chuang, 2010). Empirically, adopting the KL divergence yields larger improvements than both the Frobenius norm and the VN divergence for designing Shampoo’s schemes, as shown in Fig.˜5, and in other applications (Kulis et al., 2009).

### v3

@@ Abstract
Shampoo and its efficient variant, SOAP, employ structured second-moment estimations and have shown strong performance for training neural networks (NNs). In practice, however, Shampoo typically requires step-size grafting with Adam to be competitive, and SOAP mitigates this by applying Adam in Shampoo’s eigenbasis—at the cost of additional memory overhead from Adam in both methods. Prior analyses have largely relied on the Frobenius norm to motivate these estimation schemes. We instead recast their estimation procedures as covariance estimation under Kullback-Leibler (KL) divergence minimization, revealing a previously overlooked theoretical limitation and motivating principled redesigns. Building on this perspective, we develop KL-Shampoo and KL-SOAP, practical schemes that match or exceed the performance of Shampoo and SOAP in NN pre-training while achieving SOAP-level per-iteration runtime. Notably, KL-Shampoo does not rely on Adam to attain competitive performance, eliminating the memory overhead introduced by Adam. Across our experiments, KL-Shampoo consistently outperforms SOAP, Shampoo, and even KL-SOAP, establishing the KL-based approach as a compelling foundation for designing structured methods in NN optimization.
@@ 1 Introduction
Optimizer Shampoo (Gupta et al., 2018) has recently dethroned Adam (Kingma & Ba, 2015) as the winner of several competitions in training a wide range of neural network (NN) models (Dahl et al., 2023; Kasimbeg et al., 2025). Consequently, Shampoo and its variant, SOAP (Vyas et al., 2025a), have drawn increasing attention. In practice, Shampoo requires step-size grafting with Adam to achieve competitive performance (Agarwal et al., 2020; Shi et al., 2023). SOAP addresses this by applying Adam in Shampoo’s eigenbasis and further reducing per-iteration runtime. However, reliance on Adam introduces additional memory overhead in both methods. Prior work (Morwani et al., 2025; Eschenhagen et al., 2025; An et al., 2025; Xie et al., 2025) has investigated their structural preconditioner schemes—which approximate the flattened gradient 2nd moment (Duchi et al., 2011)—through the Frobenius norm. However, few studies have examined these schemes from the perspective of Kullback–Leibler (KL) divergence. Compared to the Frobenius norm, the KL divergence between zero-mean Gaussian covariance matrices is more appropriate for interpreting Shampoo’s and SOAP’s preconditioners as Gaussian covariance (Amari, 2016; Minh & Murino, 2017), since the second moment they approximate can be viewed as the covariance matrix of a zero-mean Gaussian. A similar KL perspective has provided a unified framework to interpret (Fletcher, 1991; Waldrip & Niven, 2016) and extend (Kanamori & Ohara, 2013a, b) structural preconditioner estimation in quasi-Newton methods such as BFGS and DFP—something the Frobenius norm does not. Moreover, the KL divergence intrinsically respects the symmetric positive-definite constraint (Amari, 2016; Minh & Murino, 2017) that preconditioners in Shampoo and SOAP must satisfy as adaptive (preconditioned) methods (Nesterov et al., 2018)—a property the Frobenius norm lacks. This constraint implies that the entries of the preconditioning matrix do not play equivalent roles and therefore should not be treated equally (Pennec et al., 2006; Bhatia, 2007)—a point the Frobenius norm ignores.
@@
In this work, we introduce a KL perspective that interprets the estimation schemes of Shampoo and SOAP as solutions to KL-minimization problems for covariance estimation. Our approach naturally extends to tensor-valued settings, where some existing theoretical interpretations may not apply. This perspective reveals a key limitation (illustrated in Fig.˜5): the Kronecker-structured estimators used by Shampoo and SOAP do not adequately solve the corresponding KL-minimization problem. This limitation, in turn, opens new opportunities for improvement. Leveraging this insight, we refine the estimation rules of Shampoo and SOAP and develop practical KL-based schemes—KL-Shampoo and KL-SOAP—that meet or exceed the performance of Shampoo and SOAP for NN (pre-)training while maintaining SOAP-level per-iteration runtime. Notably, KL-Shampoo does not rely on Adam to achieve competitive performance, thereby avoiding Adam’s additional memory overhead (Footnote˜1). In addition, our practical techniques developed for KL-Shampoo (Sec.˜4) can be adapted to strengthen existing Shampoo variants and make the trace-scaling variant competitive without step-size grafting (Fig.˜10, Appx.˜H), while also achieving SOAP-level per-iteration runtime (Fig.˜10, Appx.˜H). Empirically, we show that KL-based methods are competitive for training a range of NNs and remain as flexible as Shampoo and SOAP for tensor-valued weights. Surprisingly, KL-Shampoo consistently outperforms the other methods in our experiments (Figs.˜1 and 5). Overall, our KL-based approach provides a principal way for designing structured methods in NN optimization.
@@ 2 Background
For presentation simplicity, we focus on matrix-valued weights and the optimization update for a single parameter matrix ${\bm{\Theta}}\in\mathcal{R}^{d_{a}\times d_{b}}$, rather than a set of weight matrices for NN training. We use $\operatorname{Mat}(\cdot)$ to unflatten its input vector into a matrix and $\operatorname{vec}(\cdot)$ to flatten its input matrix into a vector. For example, ${\bm{\theta}}:=\operatorname{vec}({\bm{\Theta}})$ is the flattened weight vector and ${\bm{\Theta}}\equiv\mathrm{Mat}({\bm{\theta}})$ is the unflattened weight matrix. Vector $\bm{g}$} is a (flattened) gradient vector for the weight matrix. We denote $\gamma$, $\beta_{2}$ and $\bm{S}$} to be a step size, a weight for moving average, and a preconditioning matrix for an adaptive method, respectively. $\mathrm{Diag}(\cdot)$ returns a diagonal matrix whose diagonal entries are given by its input vector, whilst $\mathrm{diag}(\cdot)$ extracts the diagonal entries of its input matrix as a vector.
@@
Using Adam for Shampoo’s stabilization increases memory usage. If the eigendecomposition is applied infrequently to reduce iteration cost, Shampoo has to apply step-size grafting with Adam to maintain performance (Agarwal et al., 2020; Shi et al., 2023) as empirically shown in Fig.˜2. Unfortunately, this increases its memory usage introduced by Adam (see Footnote˜1).
@@
Runing Adam in the eigenbasis increases memory usage. Introducing Adam’s (RMSProp’s) 2nd moment estimation increases SOAP’s memory consumption (see Footnote˜1). This is because this estimation, ${\bm{d}}\in\mathcal{R}^{d_{a}d_{b}\times 1}$, uses extra memory and cannot be expressed as a Kronecker product of any two vectors, such as ${\bm{d}}\neq{\bm{r}}_{a}\otimes{\bm{r}}_{b}$, where ${\bm{r}}_{a}\in\mathcal{R}^{d_{a}\times 1}$ and ${\bm{r}}_{b}\in\mathcal{R}^{d_{b}\times 1}$.
@@
Table 1: Memory usage of each method considered in this work. We store and update  [^1] Since the current QR/eigen implementation in PyTorch does not support half-precision (bfloat16), we perform QR/eigen to compute ${\bm{Q}}_{k}$ for $k\in\{a,b\}$ in single precision (float32) and then cast and store them in half precision (bfloat16). Other matrices and vectors can be computed and updated in half precision. all of them in half precision (bfloat16). The memory overhead introduced by Adam is highlighted in red. Note that SOAP’s and KL-SOAP’s preconditioners, ${\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$, can not be expressed as a Kronecker product due to the augmented eigenvalues $\bm{d}$}, while Shampoo’s and KL-Shampoo’s preconditioners, ${\bm{Q}}\mathrm{Diag}({\bm{\lambda}}_{a}\otimes{\bm{\lambda}}_{b}){\bm{Q}}^{\top}$, can, where ${\bm{Q}}:={\bm{Q}}_{a}\otimes{\bm{Q}}_{b}$.
@@ 3 Second Moment Estimation via Kullback–Leibler Minimization
Many works (Morwani et al., 2025; Eschenhagen et al., 2025; An et al., 2025; Xie et al., 2025) primarily focus on matrix-valued weights and interpret Shampoo’s and SOAP’s estimation rules from the Frobenius-norm perspective. However, this norm does not account for the SPD constraint implicitly imposed on Shampoo’s and SOAP’s preconditioners, which ensures that the preconditioned gradient direction is a descent direction (Nesterov et al., 2018). As emphasized in the literature (Pennec et al., 2006; Bhatia, 2007), it is more appropriate to consider a “distance” that respects this constraint. We adopt the KL divergence because it naturally incorporates the SPD constraint, is widely used for covariance estimation (Amari, 2016; Minh & Murino, 2017), and provides a unified framework for reinterpreting and improving Shampoo’s estimation—even in tensor-valued settings where some existing interpretations based on singular value decomposition (Van Loan & Pitsianis, 1993) may not apply. In numerical optimization, the KL divergence, known as a merit function (Byrd & Nocedal, 1989), offers a unifying interpretation (Fletcher, 1991; Waldrip & Niven, 2016) and extension (Kanamori & Ohara, 2013a, b) of the low-rank estimation schemes of BFGS and DFP. In contrast, the standard Frobenius norm cannot recover these updates without additional weighting (see Sec. 6.1 of Nocedal & Wright (2006)). In statistical estimation and inference, this KL divergence is also preferred over the Frobenius norm (James et al., 1961; Kivinen & Warmuth, 1999; Davis & Dhillon, 2006; Khan & Lin, 2017; Lin et al., 2019; Kunstner et al., 2021).
@@ 3.2 Improving Shampoo’s estimation: Idealized KL-Shampoo
Our KL perspective reveals a key limitation—empirically demonstrated in Fig.˜5—of Shampoo’s Kronecker estimation with $p=\nicefrac{{1}}{{2}}$ as a one-sided approach: it does not adequately solve the KL-minimization problem when both factors are learned jointly. Motivated by this, we design an improved estimation rule that updates the two factors simultaneously. We term this scheme as idealized KL-Shampoo, which is a two-sided approach.
@@
˜2 (proof in Appx.˜B) establishes a closed-form condition (see Eq.˜4) when simultaneously learning both Kronecker factors to minimize the KL problem. Statistically, this condition corresponds to the maximum-likelihood covariance estimator of a zero-mean matrix Gaussian (Dutilleul, 1999). This condition implies that $\bm{G}$} is generated from a zero-mean matrix Gaussian with row-wise covariance ${\bm{S}}_{a}^{*}$ and column-wise covariance ${\bm{S}}_{b}^{*}$. Under this condition, Shampoo-style preconditioning naturally induces matrix-Gaussian (row and column) whitening. Our KL-based approach further extends this to tensor-Gaussian whitening for tensor-valued gradients—without the prohibitive cost typically associated with SVD-based methods. Notably, this kind of whitening arises from minimizing KL divergence, not the Frobenius norm. In machine learning, Lin et al. (2019, 2024) treated this condition as a theoretical example of a multilinear exponential-family (see Sec 5 of Lin et al. (2019)) for Kronecker-based optimization via natural gradient descent on an exponential-family manifold, while more recently, Vyas et al. (2025b) considered a similar condition motivated heuristically by gradient whitening. However, we cannot directly use this condition due to the correlated update between ${\bm{S}}_{a}$ and ${\bm{S}}_{b}$. For example, solving ${\bm{S}}_{a}^{*}$ requires knowing ${\bm{S}}_{b}^{*}$ in Eq.˜4 or vice versa. In practice, this optimal condition is often unachievable because the expectations in Eq.˜4 have to be approximated. Thus, we consider an estimated ${\bm{S}}_{k}$ to approximate ${\bm{S}}_{k}^{*}$ for $k\in\{a,b\}$ and propose an exponential moving average (EMA) scheme in practice:
@@
3.3 Comparison with Kronecker Schemes Using Alternative Divergences
Frobenius norm
@@ 3.3 Comparison with Kronecker Schemes Using Alternative Divergences
Morwani et al. (2025) consider a two-sided Shampoo variant based on the Frobenius norm and derive the optimal solution via rank-1 singular value decomposition (SVD) of the outer product $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ (Van Loan & Pitsianis, 1993). However, this optimal solution is often unattainable in practice and is computationally expensive for two reasons: (1) the expectation $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ must be approximated; and (2) performing the SVD itself is costly—yielding complexity ($O(d_{a}^{2}d_{b}^{2})$) in general even for rank-1 SVD—which is higher than the eigendecompositions with complexity ($O(d_{k}^{3})$) for $k\in\{a,b\}$ that we explicitly aim to avoid. Instead, we analyze the stationarity conditions (˜6, Appx.˜F) and derive a new variant, idealized F-Shampoo (Fig.˜6, Appx.˜F), that is structurally similar to KL-Shampoo. While a straightforward implementation of F-Shampoo performs poorly in practice, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted to improve its performance (see Fig.˜7, LABEL:{app:extra_exp}).
von Neumann divergence
@@
Another Shampoo variant, often discussed in the literature (Morwani et al., 2025; Vyas et al., 2025a; Eschenhagen et al., 2025), is Shampoo with trace scaling. Vyas et al. (2025a) established that Shampoo with trace scaling is equivalent to running Adafactor (Shazeer & Stern, 2018) in Shampoo’s eigenbasis. In contrast, KL-Shampoo is not equivalent to running Adafactor in its eigenbasis. To clarify this distinction, we make the theoretical connection between Shampoo and Adafactor more explicit: Shampoo with trace scaling is exactly a matrix generalization of Adafactor obtained by minimizing the von Neumann (VN) divergence (Tsuda et al., 2005; Dhillon & Tropp, 2008; Nock et al., 2012) and recovers Adafactor when its Kronecker factors are restricted to be diagonal, as we establish in ˜7 (Appx.˜G). By contrast, KL-Shampoo minimizes the KL divergence instead of the VN divergence. A straightforward implementation of Shampoo with trace scaling—referred to as idealized VN-Shampoo—performs poorly in practice. Howeve, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted (Fig.˜8, Appx.˜G) to substantially improve its performance, as shown in Fig.˜10 (LABEL:{app:extra_exp}).
@@
A natural question then arises: which “distance” is more suitable? Theoretically, the KL divergence is broadly applicable to arbitrary SPD matrices (Bhatia, 2007; Boumal et al., 2014) and is widely used for covariance matrices (Minh & Murino, 2017). In contrast, the Frobenius norm does not respect the SPD constraint, and the VN divergence is primarily motivated, studied, and applied for unit-trace SPD matrices (Tsuda et al., 2005; Nielsen & Chuang, 2010). Empirically, adopting the KL divergence yields larger improvements than both the Frobenius norm and the VN divergence for designing Shampoo’s schemes (see Fig.˜5) and in other applications (Kulis et al., 2009).